# distributed manipulator systems DoF degrees of freedom

### v1

@@ IV-C Workspace - Connected Tiles
Figure 4(a) illustrates how $\alpha$ varies as $\delta$ and $\phi$ change for two connected tiles at a fixed $r_{0}$.The plot reveals $\delta$ and $\phi$ have a complex, non-linear effect on $\alpha$.
@@ VII Discussion
We demonstrated translation with simple geometric objects. While this serves as a proof-of-concept, objects with different geometries may not follow similar trajectories and could require modified movement patterns for reliable translation. Unconventional geometries may also move in a less predicable manner, necessitating new strategies for reliable manipulation. Although our focus is on objects smaller than the inter-actuator spacing, larger objects that contact multiple tiles simultaneously would require different manipulation strategies, such as gating. We would also like to investigate the manipulation of deformable objects, as the lack of grasping may be advantageous for such objects.
Our experiments used an open-loop, hand-crafted manipulation strategy to demonstrate the system’s functionality and the role of the connective material. Future work should explore alternative tile movement strategies, including closed-loop control for more precise object manipulation capabilities. This opens the possibility for more precise control and for more precise tile movement to induce particular object motion, as its current location is known. It also means modifications to movement, such as vibration, can be selectively applied only in cases where the object becomes stuck.
Here we consider the material as inextensible. This is not true of real-world materials, though for the object mass and motor torques investigated, this approximation holds for our use case. In the real world, compliance may cause hysteresis and energy loss, potentially reducing manipulation precision and repeatability. Future work should examine how material stretch affects object manipulation, reachable workspaces, positional drift over cycles, and motor loading, as well as exploring ways to exploit or mitigate this behaviour.
@@ VIII Conclusion
This work introduces a soft, interconnected actuator array that enables effective object manipulation with significantly reduced actuator density. By interconnecting actuators with a flexible material the proposed design not only extends the reachable workspace but also relaxes traditional scale and density constraints inherent to distributed manipulator systems. Through kinematic analysis and the implementation of a sinusoidal and a cycling state based movement patterns, the system successfully translated objects of different geometries across a large distance, achieving a $2.51\times$ increase in traversable distance when compared to a similar, densely packed system that did not use connective material.

### v2

@@ IV-C Workspace - Connected Tiles
Figure˜4(a) illustrates how $\alpha$ varies as $\delta$ and $\phi$ change for two connected tiles at a fixed $r_{0}$.The plot reveals $\delta$ and $\phi$ have a complex, non-linear effect on $\alpha$.
@@ VII Discussion
We demonstrated translation with simple geometric objects. While this serves as a proof-of-concept, objects of other geometries may follow different trajectories and require modified movement patterns for reliable translation. Irregular geometries may also move in a less predictable manner, necessitating novel manipulation strategies. Although our focus is on objects smaller than the inter-actuator spacing, larger objects that contact multiple tiles simultaneously would require different manipulation strategies, such as gating. We would also like to investigate the manipulation of deformable objects, for which the absence of grasping may be advantageous.
Our experiments used an open-loop, hand-crafted manipulation strategy to demonstrate the system’s functionality and the role of the connective material. Future work should explore closed-loop control and manipulation strategies for greater manipulation precision, enabling targeted motions based on object location and allowing selective application of movements such as vibration when an object becomes stuck.
Here we consider the material as inextensible. This is not true of real-world materials; though this approximation holds for the object masses and motor torques tested. In the real world, compliance may cause hysteresis and energy loss, potentially reducing manipulation precision and repeatability. Future work should examine how material stretch affects object manipulation, reachable workspaces, positional drift over cycles, and motor loading, as well as exploring ways to exploit or mitigate this behaviour.
@@ VIII Conclusion
This work introduces a soft, interconnected actuator array that enables effective object manipulation with significantly reduced actuator density. By interconnecting actuators with a flexible material the proposed design not only extends the reachable workspace but also relaxes traditional scale and density constraints inherent to distributed manipulator systems. Through kinematic analysis and the implementation of a sinusoidal and a cycling state based movement patterns, the system successfully translated objects of different geometries across a large distance, achieving a $1.84\times$ increase in .traversable distance when compared to a similar, densely packed system that did not use connective material.